OpenStudy (anonymous):

multiply the write answer without the exponents (12x10^5)x(3x10^6)

jimthompson5910 (jim_thompson5910):

12 * 3 = ???

OpenStudy (anonymous):

i have a answer but idk if it is write

jimthompson5910 (jim_thompson5910):

ok show me

OpenStudy (anonymous):

2 multiply by 3=6 10^5 and 10^6= 10^1 so 6x10^1=0.06

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

what is 12 times 3 ?

OpenStudy (anonymous):

36

jimthompson5910 (jim_thompson5910):

What is 10^5 times 10^6 ?

OpenStudy (anonymous):

im sorry its (2x10^5)

OpenStudy (anonymous):

i typed too fast

jimthompson5910 (jim_thompson5910):

rewrite the whole problem again

OpenStudy (anonymous):

ok (2x10^5) x (3x10^6)

jimthompson5910 (jim_thompson5910):

so 2 * 3 = 6, good so far

jimthompson5910 (jim_thompson5910):

10^5 times 10^6 = ???

OpenStudy (anonymous):

10^1

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

5-6

jimthompson5910 (jim_thompson5910):

you add the exponents

jimthompson5910 (jim_thompson5910):

5+6 = 11, so 10^5 times 10^6 = 10^11

OpenStudy (anonymous):

6x10^11=

jimthompson5910 (jim_thompson5910):

so the product (2x10^5) x (3x10^6) becomes 6 x 10^11

jimthompson5910 (jim_thompson5910):

now write that in standard form

jimthompson5910 (jim_thompson5910):

basically, you would write 11 zeros after the '6'

OpenStudy (anonymous):

60^11

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

6 x 10^11 = 600,000,000,000 This is the number 600 billion

OpenStudy (anonymous):

so i had a similar problem to that and i had to subtract the exponents

jimthompson5910 (jim_thompson5910):

when you divide, yes, you subtract the exponents

jimthompson5910 (jim_thompson5910):

but this time you're multiplying, so you add the exponents

OpenStudy (anonymous):

oh thats y

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

and i kept putting 0.06... ok i have another question

jimthompson5910 (jim_thompson5910):

ok you can ask one more

OpenStudy (anonymous):

simplify (6a)^-2

OpenStudy (anonymous):

i was wondering do i multiply 6x6=

jimthompson5910 (jim_thompson5910):

yes, that's one of the steps, but not yet though

jimthompson5910 (jim_thompson5910):

when you have a negative exponent, you can use this idea x^(-k) = 1/(x^k) where k is any positive number

jimthompson5910 (jim_thompson5910):

so for example, x^(-2) = 1/(x^2)

jimthompson5910 (jim_thompson5910):

this means that (6a)^(-2) = 1/( (6a)^2 )

OpenStudy (anonymous):

ok im viewing

jimthompson5910 (jim_thompson5910):

now square 6a, what do you get

OpenStudy (anonymous):

ok i get 36a

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

a*a = ???

OpenStudy (anonymous):

im not sure 6x6=36

jimthompson5910 (jim_thompson5910):

good, so a times a = a^2

jimthompson5910 (jim_thompson5910):

which means (6a)^2 = 36a^2

jimthompson5910 (jim_thompson5910):

and that (6a)^(-2) = 1/( 36a^2 )

OpenStudy (anonymous):

so 36a^2 as the answer

jimthompson5910 (jim_thompson5910):

1 over 36a^2

jimthompson5910 (jim_thompson5910):

36a^2 is the entire denominator

OpenStudy (anonymous):

1/36a^2

jimthompson5910 (jim_thompson5910):

yes \[\Large \frac{1}{36a^2}\]

OpenStudy (anonymous):

oh ok well its another similar to that but different

jimthompson5910 (jim_thompson5910):

i gotcha

OpenStudy (anonymous):

it says (-6p)^4

jimthompson5910 (jim_thompson5910):

that's similar, but there's no negative exponent

jimthompson5910 (jim_thompson5910):

so you won't have a fraction

OpenStudy (anonymous):

i have 1296p

jimthompson5910 (jim_thompson5910):

it should be 1296p^4

jimthompson5910 (jim_thompson5910):

since you also raise p to the 4th power

OpenStudy (anonymous):

oh ell you c im trying

OpenStudy (anonymous):

i thank u

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

do you have time for one more to see if im correct?

OpenStudy (anonymous):

hello